OpenStudy (anonymous):

find f o g if f(x) = x+4 and g(x) = 2x -1

OpenStudy (anonymous):

@Michele_Laino @AlexandervonHumboldt2 @Luigi0210 @nincompoop

OpenStudy (danjs):

f [ g(x) ]

OpenStudy (danjs):

chained together compound function, instead of f(x), with x as the input, you have f[ g(x)] with the whole g(x) as the input

OpenStudy (anonymous):

WHich gives me what

OpenStudy (danjs):

f(x) = x + 4 instead of x, use g(x) as input for the compound function f[g(x)] = g(x) + 4 = (2x - 1) + 4 that is it, simplify it if you want

OpenStudy (bluemiku):

lol i was gonna say something similar to that @DanJS

OpenStudy (danjs):

f(@BlueMiku ) = @BlueMiku + 4